OpenStudy (anonymous):

MEDAL!! If h(x)=(fog)(x) and h(x)=3(x+2)^2, find f(x) and g(x) A. f(x)=3x^2 & g(x)=(x+2)^2 B. f(x)=(x+2)^2 & g(x)=3x^2 C. f(x)=x+2 & g(x)=3x^2 D. f(x)=3x^2 & g(x)=x+2

OpenStudy (praxer):

i suppose it is D because if we take g(x)=y=x+2 than f(y)=3y^2 =3(x+2)^2 =h(x) =fog(x)